# TASER: Table Agents for Schema‑guided Extraction and Recommendation

### v1

@@ Abstract
Real-world financial documents report essential information about an entity’s financial holdings that can span millions of different financial instrument types. Yet, these details are often buried in messy, multi-page, fragmented tables - for example, 99.4% of the tables in our dataset have no bounding boxes with the maximum number of rows amounting to 426 per table across 44 pages. To tackle these unique challenges from real-world tables, we present a continuously learning, agentic table extraction system, TASER (Table Agents for Schema-guided Extraction and Recommendation) that extracts highly unstructured, multi-page, heterogeneous tables into normalized, schema-conforming outputs. Our table agents execute on table detection, classification, extraction, and recommendations by leveraging an initial schema. Then, our Recommender Agent reviews the outputs, recommends schema revisions, and decides on the final recommendations, enabling TASER to outperform existing table detection models such as Table Transformer by 10.1%. Within this continuous learning process, we highlight that larger batch sizes result in a 104.3% increase in schema recommendations that are actionable and utilized, resulting in a 9.8% increase in extracted holdings - highlighting the importance of a continuous learning process. To train TASER, we have manually labeled 22,584 pages (28,150,449 tokens), 3,213 tables for \$731,685,511,687 of holdings culminating in one of the first real financial table datasets. We release our dataset TASERTab to enable the research community to access real-world financial tables and outputs. Our results highlight the promise of agentic, schema-guided extraction systems for robust understanding of real-world financial tables.
Figure 1: Complexity of Holdings Table in Regulatory Filings. In the original format, multiple data attributes are displayed in a single line, with no bounding boxes, rendering the generation of structured outputs highly challenging. TASER enables the generation of structured outputs from highly variable, multi-page financial tables for complex instrument holdings. Negative quantities or market values denote short positions. See Appendix L for additional outputs.
Figure 2: Variety and complexity of financial tables. From leftmost column - for a single financial filing type, such as annual reports, there are a variety of formats by which these reports are filed. Within each filing, there are numerous table types - spanning from Performance, Financial Holdings, or Cash Flow Tables. Each table taxonomy houses very different types of information and has different objectives. For example, the Cash Flow Table houses information relating to operational cash movements for an entity. In contrast, the Financial Holdings Table (which is our table of interest in this paper) displays all financial instruments that an entity holds. For the Financial Holdings Table, there are numerous layout structures and formats by which tables appear, as seen in the rightmost column. Due to the extreme heterogeneity of formatting, document layout, and table structure, traditional table extraction methods fail to perform for financial filings.
Model Modality Primary Task Promptable
Camelot Vision + Spatial Heuristic Table Detection & Parsing No
Table Transformer Vision Detection & Structure Recognition No
TaPas Text Table‑based QA Partially
TAPEX Text Programmatic Extraction (SQL) Partially
TASER (ours) Vision + Text Schema‑guided Extraction Yes
Table 1: Comparison of representative table extraction and reasoning models. Our work extends prior methods by introducing a fully agentic, schema-guided extraction framework for highly complex financial tables, leveraging prompt-based self-refinement and continuous schema adaptation.
@@ 1 Introduction
Financial documents, particularly annual regulatory filings for funds, house tables that govern \$68.9 trillion of investments globally (Investment Company Institute 2024). In perspective, \$68.9 trillion is more than twice the total Gross Domestic Product (GDP) of the United States (\$29.1 trillion) (WorldBank 2025). This critical data is housed in the Financial Holdings Table (Figure 1), which outlines the entirety of an entity’s investment holdings (U.S. Congress 1934; EU Commission 2019); this table is the longest table in terms of row count (with a maximum row count of 426 rows) - more than double the average row count of all other table types (Table 7). These Financial Holdings Tables are not only very long but also highly heterogeneous in layout structure and visually complex (Figure 2). While generating structured outputs from these tables is critical for many regulatory and financial institutions (Cho et al. 2024), there is a relative dearth of studies that focus on understanding the complexity of Financial Holdings Tables, compared to web or SQL tables (Herzig et al. 2020; Pasupat and Liang 2015; Zhong, Xiong, and Socher 2017). Therefore, the following challenges exist in terms of parsing Financial Holding Tables into structured, machine-readable outputs (Figure 2):
[…]
    One-to-many relationships between a document and the tables it houses exacerbates standard model performance for table detection or structure recognition tasks.
[…]
    Financial Holdings Tables span across multiple pages, rendering models that operate at the page-level inefficient.
[…]
    Financial instruments are highly complex, with embedded structures and nested hierarchies. Therefore, details are often clumped in a single cell as seen in Figure 1.
[…]
    Tabular layouts are heterogeneous with no bounding boxes, mixing tables, text blocks, footnotes, and images, often without consistent labeling or alignment. Moreover, we identify that 99.4% of tables in our dataset, TASERTab, lack bounding boxes or gridlines to efficiently identify a single cell.
These challenges drive our motivation to construct our agentic table extraction methodology that is capable of goal-driven parsing and self-refinement, with the ability to decide on the optimal schema, continuously learning and reasoning from errors.
@@
Contribution 1: We propose a novel table extraction methodology, TASER (Table Agents for Structured Extractions and Recommendation) that undertakes table detection, classification, extraction, and recommendations in a singular pipeline by leveraging the provided schema. We compare our methodology against predominant table extraction methodologies and report TASER’s 10.1% outperformance against Table Transformer (Yang et al. 2022) for detection.
@@
Contribution 2: Once the initial schema is given, we prove the effectiveness of our Recommender Agent to continuously improve the initial schema - reflecting a tunable and continuous self-learning loop. Throughout our training, we discover that small batches are optimal in providing diverse and comprehensive recommendations to the original schema–however, at the cost of redundant recommendations. In contrast, large batches drive high precision recommendations - however, at the cost of diversity. Thus, our results establish that self-learning via agents for table extraction is tunable; through adjusting batch size, we can directly control schema refinement to maximize actionable coverage while minimizing redundancy.
@@
Contribution 3: We have constructed a manually labeled dataset TASERTab of ground truth labels for 3213 real-world Financial Holdings Tables amounting to \$731.7B in value. We sourced the filings directly from fund websites, labeled the total net assets for each fund, and recorded the span of each Financial Holdings Table. We believe that this is the first dataset of its kind for the research community to access real-world financial tables with structured outputs.
@@ 2 Related Work
Information & Table Extraction: Early information extraction relied on statistical models (HMMs (Borkar, Deshmukh, and Sarawagi 2001), CRFs (Lafferty, McCallum, and Pereira 2001), heuristics (Press 2003), and graph-based layouts (Liu et al. 2019; Qian et al. 2019; Meuschke et al. 2023), but still struggle with complex, heterogeneous tables.
Table Representation Learning: Transformer-based table understanding and QA include TaPaS (Herzig et al. 2020), TaBERT (Yin et al. 2020), TaPEX (Liu et al. 2022), TURL (Deng et al. 2020), TUTA (Wang et al. 2021), and TableFormer (Yang et al. 2022). These methods encode text, structure, and layout, but few are benchmarked on long, dense, multi-page financial reports.
LLMs for Structured Data: General LLMs have strong performance for schema-conformant extraction via fine-tuning & prompting (Brown et al. 2020; Ouyang et al. 2022; Paolini et al. 2021; Wei et al. 2023; Lu et al. 2022; Liu and Contributors 2024), while multimodal approaches (LayoutLM (Xu et al. 2020, 2021; Huang et al. 2022), DONUT (Kim et al. 2022), DocFormer (Appalaraju et al. 2021), UniTable (Peng et al. 2024), and Table Transformer (Smock, Pesala, and Abraham 2021; Carion et al. 2020) improve layout awareness but still lag on long, fragmented tables (Zhao et al. 2024).
Financial Document Parsing: Includes multimodal extraction (Watson and Liu 2020), expert agent pipelines (Cho et al. 2024), and large-scale benchmarks (DocILE (Šimsa et al. 2023), BuDDIE (Wang et al. 2025), Kleister (Stanisławek et al. 2021), FinTabNet (Zheng et al. 2020)).
Agentic and Recursive Extraction: Recent methods cast LLMs as agents capable of iterative extraction and self-correction (Shen et al. 2023; Roucher et al. 2025; Yao et al. 2023; Shinn et al. 2023; Watson et al. 2023; Yuan and Xie 2025). Prompt-based feedback, introspective refinement, and episodic memory frameworks (Madaan et al. 2023; Shinn et al. 2023; Yao et al. 2023; Wei et al. 2023; Wang et al. 2023; Shridhar et al. 2023; Du et al. 2024; Lee et al. 2024; Lewis et al. 2021; Schulhoff et al. 2025) drive improvements in reasoning for complex extraction.
LLM-Augmented Aggregation: LLMs are increasingly used for hierarchical and agglomerative clustering, enhancing interpretability and item-set quality (Viswanathan et al. 2024; Pattnaik et al. 2024; Zhang, Wang, and Shang 2023) by leveraging semantic reasoning in merge decisions.
@@ 3.1 System Architecture
    Detector Agent: Identifies candidate pages containing Financial Holdings Tables leveraging the initial schema provided. The prompt is tuned to maximize recall as we do not want to miss any Financial Holdings Tables. We provide our prompts in the Appendix (Figure 12).
@@
    Recommender Agent: Reviews unmatched extractions containing both false and true positives. A false positive is a spurious extraction (e.g., headers/subtotals/footnotes/OCR noise or cells from non-holdings tables) that fails schema/type/consistency checks; a true positive is a valid holdings field from a genuine row that the current schema cannot yet classify (e.g., new security type or column alias) but passes those checks. The agent first filters false positives by re-validating each candidate under the current schema; it proposes schema modifications for the remaining true positives and, per class, recommends the minimal change that correctly classifies each item.
All agents interact through explicit artifacts: structured outputs, episodic error stacks, and schema definitions. Output validation is integrated into each agent’s forward pass via instructor (Liu and Contributors 2024). TASER implements a recursive feedback loop, where errors and unmatched holdings identified in the initial extraction are escalated to the Recommender Agent, which provides recommendations to refine the schema and triggers re-extraction. This loop repeats until all entries are either matched or a stopping criterion is met (no new schema modifications proposed). The pipeline is fully parallelizable for multi-page and multi-entity filings. A schematic of the full agentic pipeline is shown in Figure 3.
Figure 3: Agentic feedback loop for schema-guided table extraction and continuous schema recommendations. The pipeline begins with raw document ingestion, where candidate pages are evaluated by a Table Detector Agent. Only pages identified as Financial Holdings Tables proceed to the extraction stage, where our Extractor Agent parses each table into a structured, schema-conformant response (e.g., extracting option instrument fields as shown). In the Refinement stage, outputs are partitioned into declared (matched) and unmatched holdings. Unmatched holdings are grouped into batches, which trigger a schema recommendation and feedback loop from the Recommender Agent (Figure 4). The Recommender Agent proposes schema refinements, reconciles its recommendations, and finally, updates the schema, enabling TASER to adapt and improve extraction fidelity over successive iterations.
Figure 4: Schema-Guided Agentic Extraction and Refinement Loop. The extraction pipeline begins with an Initial Schema Definition (v1), which guides the LLM Extractor Agent as it processes the raw Holdings Table to produce Declared Holdings. Holdings that do not match the schema are routed as Unmatched Holdings, triggering the generation of Schema Update Suggestions. These suggestions are reviewed, clustered, and aggregated by our Validation Agent before updating the schema (v2), replacing the prior definition and closing the agentic feedback loop. This process enables continuous schema refinement and robust extraction from heterogeneous, visually complex financial tables.
@@ 3.2 Initial Schema Definition and Application
TASER’s extraction process is anchored by an explicit, user-modifiable Portfolio schema that defines the target structure for Financial Holdings Tables. We implement this schema using Pydantic models; our initial schema reflects our current knowledge by leveraging external knowledge (U.S. Congress 1934). Each schema consists of a base Instrument model, subclassed for common asset types (e.g., Equity, Bond, Option, Swap, Forward, Future, Debt, EquityLinkedNote) and an Other class for uncategorized assets. Each subclass specifies instrument-specific fields and validation logic (see App. H).
Schema-Guided Extraction: For each candidate page, the Extractor Agent prompts the LLM with the current schema embedded in the prompt context. The LLM is instructed to return a structured output, which is immediately parsed and validated against the schema using Pydantic’s type checking and validation logic. Outputs that fail schema validation (e.g., missing fields, type errors, or undeclared instruments) are flagged as feedback for re-extraction.
Schema Recommendations for Iterative Refinement: We formalize schema refinement as an iterative, LLM-driven clustering process that updates the original schema to accommodate unmatched holdings and novel instrument types discovered during extraction (Novikov et al. 2025; Pattnaik et al. 2024; Zhang, Wang, and Shang 2023). At each iteration, the agent operates on the episodic error stack to propose schema modifications, and extraction is retried using the updated schema. This process continues until all entries are matched or no further improvements are possible. Let $H=\{h_{1},h_{2},\ldots,h_{N}\}$ denote the set of unmatched holdings, and let $\Sigma^{(0)}$ be the initial schema. For each iteration $\ell$, we define:
@@
    Aggregate all proposals, cluster/reconcile if necessary, and select the modifications to apply.
@@ 3.3 Ablation Strategies and Efficiency
We conduct a systematic ablation study to isolate the contributions of schema-guided extraction, prompt engineering, and agentic feedback within TASER. Four main extraction strategies are compared:
@@
Table 2 reports detection and extraction accuracy for each strategy. Full Schema Prompting achieves the highest precision, F1, and accuracy, while maintaining perfect recall. Table 3 summarizes extraction fidelity in terms of total absolute dollar difference and percent unaccounted value, and Table 10 compares computational efficiency in terms of tokens and latency.
Method Recall (%) Precision (%) F1 (%) Accuracy (%)
Camelot (Hybrid) 56.92 23.46 33.23 47.35
Table Transformer (Smock, Pesala, and Abraham 2021) 99.76 32.75 49.31 46.27
TASER
(a) Raw Text Prompting 100.0 38.62 55.73 58.38
(b) Structured CoT 100.0 34.42 51.21 50.10
(c) Full Schema Prompting 100.0 43.43 59.44 66.35
(d) Direct Schema Application 100.0 41.84 58.30 63.99
Table 2: Detection performance across ablation strategies. While all TASER ablations achieve perfect recall (100%), Full Schema Prompting yields the highest precision (43.43%), F1 score (59.44%), and overall accuracy (66.35%), underscoring the value of embedding the complete Portfolio schema in the detection prompt. See Table 4 for all Camelot configurations.
Method Total Absolute Difference (USD) Unaccounted (% of total)
TASER
(a) Raw Text Prompting \$ 107,066,845 0.015%
(b) Structured CoT \$ 120,577,458 0.016%
(c) Full Schema Prompting \$ 102,836,797 0.014%
(d) Direct Schema Application \$ 118,881,312 0.016%
Table 3: Extraction metrics for each ablation strategy. We report the total absolute difference (TAD) between extracted and ground‐truth net asset values and the percentage of total holdings unaccounted for out of \$731.7 billion. Lower TAD and unaccounted percentages indicate higher dollar‐value fidelity; Full Schema Prompting achieves the best performance (TAD = \$102.8M, 0.014% unaccounted), demonstrating superior accuracy in financial extraction.

### v4

@@ Abstract
Real-world financial filings report critical information about an entity’s investment holdings, essential for assessing that entity’s risk, profitability, and relationship profile. Yet, these details are often buried in messy, multi-page, fragmented tables that are difficult to parse, hindering downstream QA and data normalization. Specifically, 99.4% of the tables in our financial table dataset lack bounding boxes, with the largest table spanning 44 pages. To address this, we present TASER (Table Agents for Schema-guided Extraction and Recommendation), a continuously learning, agentic table extraction system that converts highly unstructured, multi-page, heterogeneous tables into normalized, schema-conforming outputs. Guided by an initial portfolio schema, TASER executes table detection, classification, extraction, and recommendations in a single pipeline. Our Recommender Agent reviews unmatched outputs and proposes schema revisions, enabling TASER to outperform vision-based table detection models such as Table Transformer by 10.1%. Within this continuous learning process, larger batch sizes yield a 104.3% increase in useful schema recommendations and a 9.8% increase in total extractions. To train TASER, we manually labeled 22,584 pages and 3,213 tables covering \$731.7 billion in holdings, culminating in TASERTab to facilitate research on real-world financial tables and structured outputs. Our results highlight the promise of continuously learning agents for robust extractions from complex tabular data.
TASER: Table Agents for Schema‑guided Extraction and Recommendation
Nicole Cho    Kirsty Fielding    William Watson Sumitra Ganesh   Manuela Veloso J.P. Morgan AI Research nicole.cho@jpmorgan.com
@@ 1 Introduction
Figure 1: Complexity of Holdings Table in Regulatory Filings. In the original format, multiple data attributes are displayed in a single line, with no bounding boxes, rendering the generation of structured outputs highly challenging. TASER enables the generation of structured outputs from highly variable, multi-page financial tables for complex instrument holdings. Negative quantities or market values denote short positions. See Appendix K for additional outputs.
Figure 2: Variety and complexity of financial tables. From leftmost column - for a single financial filing type, such as annual reports, there is no consistency among reports. Within each report, there are numerous table types with each type housing very different types of information. Even within a single table type (such as the Financial Holdings Table), there are numerous layout structures, as seen in the rightmost column. Due to the extreme heterogeneity of formatting, document layout, and table structure, traditional table extraction methods fail to perform for financial filings.
@@
Financial documents, particularly annual regulatory filings for funds, house tables that govern \$68.9 trillion of investments globally (Investment Company Institute, 2024). By comparison, \$68.9 trillion is more than twice the total Gross Domestic Product (GDP) of the United States (\$29.1 trillion) (WorldBank, 2025). This critical data is housed in the Financial Holdings Table (Figure 1), which outlines the entirety of an entity’s investment holdings (U.S. Congress, 1934; EU Commission, 2019); this table has the highest row count (maximum 426 rows)–more than double the average row count of all other table types (Table 7). These Financial Holdings Tables are long and highly heterogeneous in layout (Figure 2). While generating structured outputs from these tables is critical for many regulatory and financial institutions to undertake basic QA (Question-Answering) tasks using an LLM (Large Language Model) or libraries such as pandas (Cho et al., 2024), there is a relative dearth of studies that focus on continual learning to extract from Financial Holdings Tables, compared to web or SQL tables (Herzig et al., 2020; Pasupat and Liang, 2015; Zhong et al., 2017). Therefore, the following challenges exist in terms of parsing Financial Holdings Tables into structured, machine-readable outputs: (1) One-to-many relationships between a document and the tables it houses exacerbate standard model performance for table detection or structure recognition tasks. (2) Financial Holdings Tables span across multiple pages, rendering models that operate at the page level inefficient. (3) Financial instruments are highly complex with nested hierarchies. Therefore, details are often clumped in a single cell as seen in Figure 1. (4) Tabular layouts are heterogeneous with no bounding boxes, mixing tables, text blocks, footnotes, and images, often without consistent labeling or alignment. 99.4% of tables in our dataset, TASERTab, lack bounding boxes to efficiently identify a single cell. These challenges motivate our agentic table extraction methodology capable of goal-driven parsing and self-refinement, continuously learning and reasoning from errors.
@@
Contribution 1: We propose a continuously learning, agentic table extraction methodology, TASER (Table Agents for Schema-guided Extraction and Recommendation) that performs detection, classification, extraction, and recommendation in a single pipeline by leveraging the schema invoked as a tool call. TASER is layout-agnostic and can operate for tables of any format. We compare our methodology against predominant methodologies and report TASER’s 10.1% improvement over Table Transformer (Yang et al., 2022) for detection.
@@
Contribution 2: We demonstrate the effectiveness of our Recommender Agent to continuously improve the initial schema - reflecting a tunable and continuous self-learning loop. Throughout our training, we found that small batches are optimal for providing diverse and comprehensive recommendations to the original schema–however, at the cost of redundant recommendations. In contrast, large batches drive high precision recommendations at the cost of diversity. Thus, our results establish that self-learning via agents for table extraction is tunable; through adjusting batch size, we can control schema refinement to maximize actionable coverage while minimizing redundancy.
@@
Contribution 3: We have constructed a manually labeled dataset TASERTab of ground truth labels for 3,213 real-world Financial Holdings Tables amounting to \$731.7B in value. We sourced the filings from fund websites, labeled the total net assets for each fund, and recorded the span of each Financial Holdings Table. We believe that this is the first dataset of its kind to provide access to real-world financial tables side by side with structured outputs.
@@ 2 Related Work
Information & Table Extraction: Early information extraction relied on statistical models (HMMs (Borkar et al., 2001), CRFs (Lafferty et al., 2001), heuristics (Press, 2003), and graph-based layouts (Liu et al., 2019; Qian et al., 2019; Meuschke et al., 2023), but still struggle with complex, heterogeneous tables.
Table Representation Learning: Transformer-based table understanding and QA include TaPaS (Herzig et al., 2020), TaBERT (Yin et al., 2020), TaPEX (Liu et al., 2022), TURL (Deng et al., 2020), TUTA (Wang et al., 2021), and TableFormer (Yang et al., 2022). These methods encode text, structure, and layout, but few are benchmarked on long, dense, multi-page financial reports.
LLMs for Structured Data: General LLMs have strong performance for schema-conformant extraction via fine-tuning & prompting (Brown et al., 2020; Liu and Contributors, 2024), while multimodal approaches (LayoutLM (Xu et al., 2020, 2021; Huang et al., 2022), DONUT (Kim et al., 2022), DocFormer (Appalaraju et al., 2021), UniTable (Peng et al., 2024), and Table Transformer (Smock et al., 2021; Carion et al., 2020) improve layout awareness but still lag on long, fragmented tables (Zhao et al., 2024).
Financial Document Parsing: (Watson and Liu, 2020) has focused on table extraction from images while (Cho et al., 2024) has focused on expert agent pipelines. Large-scale benchmarks such as DocILE (Šimsa et al., 2023), BuDDIE (Wang et al., 2025) have also focused on financial documents.
Agentic and Recursive Extraction: Recent methods cast LLMs as agents capable of iterative extraction and self-correction (Shen et al., 2023; Roucher et al., 2025; Watson et al., 2023; Yuan and Xie, 2025). Prompt-based feedback, introspective refinement, and episodic memory frameworks (Madaan et al., 2023; Shinn et al., 2023; Yao et al., 2023) drive improvements in reasoning for complex extraction.
@@ 3.1 System Architecture
    Detector Agent: Identifies candidate pages containing Financial Holdings Tables leveraging the initial schema provided. The prompt is tuned to maximize recall to avoid missing any Financial Holdings Tables. We provide our prompts in the Appendix (Figure 12).
@@
    Recommender Agent: Reviews unmatched extractions containing both false and true positives. A false positive is a spurious extraction (e.g., headers/subtotals/footnotes/OCR noise or cells from non-holdings tables) that fails schema/type/consistency checks; a true positive is a valid holdings field from a genuine row that the current schema cannot yet classify but passes those checks. The agent first filters false positives by re-validating each candidate under the current schema; it proposes schema modifications for the remaining true positives and, per class, recommends the minimal change needed.
All agents interact through explicit artifacts: structured outputs, episodic error stacks, and schema definitions. Output validation is integrated into each agent’s forward pass via Instructor (Liu and Contributors, 2024). TASER implements a recursive feedback loop, where errors and unmatched holdings identified in the initial extraction are escalated to the Recommender Agent, which provides recommendations to refine the schema and triggers re-extraction. This loop repeats until all entries are matched. A schematic of the full agentic pipeline is shown in Figure 3.
Figure 3: TASER deployment architecture. Users submit extraction requests through a UI or API hosted in an ECS container, which enqueues an extraction task to SQS. An AWS Lambda function orchestrates the Ingest, Detection, Extraction, and Recommender Agents, persisting task status and schema cache in DynamoDB and writing intermediary and final outputs to S3; failed tasks are routed to a dead-letter queue (DLQ) for later inspection. The Recommender Agent analyzes the intermediary output file and associated error stack to propose schema enhancement recommendations. Users can accept or reject these recommendations; accepted updates are written back to the schema cache and automatically retrigger the extraction pipeline, enabling TASER to continuously refine its extraction artifacts over time.
@@ 3.2 Initial Schema Definition and Application
TASER’s extraction process is anchored by an explicit, user-modifiable Portfolio schema that defines the target structure for Holdings Tables. We implement this schema using Pydantic models; our initial schema reflects is informed by leveraging external knowledge (U.S. Congress, 1934). Each schema consists of a base Instrument model, subclassed for common asset types (e.g., Equity, Bond, Option, Swap, Forward, Future, Debt, and an Other class for uncategorized rows. Each subclass specifies instrument-specific fields and validation logic (see App. G).
Schema-Guided Extraction: For each candidate page, the Extractor Agent prompts the LLM with the current schema embedded in the prompt context. The LLM is instructed to return a structured output, which is immediately parsed and validated against the schema using Pydantic’s type checking and validation logic. Outputs that fail schema validation (e.g., missing fields, type errors, or undeclared instruments) are flagged.
Schema Recommendations for Iterative Refinement: We formalize schema refinement as an iterative, LLM-driven clustering process that updates the schema to accommodate unmatched or novel holdings discovered during extraction (Novikov et al., 2025; Zhang et al., 2023). At each iteration, the agent operates on the episodic error stack to propose schema modifications, and extraction is retried using the updated schema. This process continues until all entries are matched or no further improvements are possible. Let $H=\{h_{1},h_{2},\ldots,h_{N}\}$ denote the set of unmatched holdings, and let $\Sigma^{(0)}$ be the initial schema. For each iteration $\ell$:
@@
    Aggregate, cluster, and select recommendations
@@
Figure 4: Schema-Guided Agentic Refinement Loop. The extraction pipeline begins with an Initial Schema Definition (v1), which guides the LLM Extractor Agent as it processes the raw Holdings Table to produce Declared Holdings. Holdings that do not match the schema are routed as Unmatched Holdings, triggering the generation of Schema Update Suggestions. These suggestions are reviewed, clustered, and aggregated by our Recommender Agent before updating the schema (v2), replacing the prior definition and closing the agentic feedback loop. This process enables continuous schema refinement and robust extraction.
@@ 3.3 Ablation Strategies and Efficiency
We systematically ablate TASER to isolate the impact of schema-guided extraction, prompt engineering, and agentic feedback across four strategies:
@@
Table 2 reports detection and extraction via absolute dollar difference, and Table 10 compares computational efficiency in tokens and latency.
Provider Model Recall (%) Precision (%) F1 Score (%) Accuracy (%)
Camelot Hybrid 56.92 23.46 33.23 47.35
Microsoft Table Transformer 99.76 32.75 49.31 46.27
OpenAI gpt-4o-2024-11-20 100.00 43.43 59.44 66.35
OpenAI gpt-5-mini-2025-08-07 94.92 54.15 68.96 80.33
OpenAI gpt-4.1-2025-04-14 95.80 54.32 69.33 80.49
OpenAI gpt-5-nano-2025-08-07 95.97 55.63 70.44 81.46
OpenAI gpt-5-2025-08-07 95.97 68.16 79.71 88.75
Anthropic claude_sonnet-3-7 88.97 57.34 69.73 82.22
Amazon nova_pro-v1-0 85.90 69.45 76.84 88.07
Table 1: Detector performance across models on TASERTab. Recall, precision, F1, and accuracy are reported for baselines and the Detector Agent instantiated with different LLMs. gpt-4o-2024-11-20 attains perfect recall, while gpt-5-2025-08-07 achieves the best overall F1 and accuracy. nova_pro-v1-0 delivers the highest precision but at the cost of lower recall, illustrating the trade-off between missing holdings tables and avoiding false positives.
Method Recall (%) Precision (%) F1 (%) Accuracy (%) TAD (USD) Unaccounted
TASER (a) Raw Text Prompting 100.00 38.62 55.73 58.38 \$ 107,066,845 0.015%
(b) Structured CoT 100.00 34.42 51.21 50.10 \$ 120,577,458 0.016%
(c) Full Schema Prompting 100.00 43.43 59.44 66.35 \$ 102,836,797 0.014%
(d) Direct Schema Application 100.00 41.84 58.30 63.99 \$ 118,881,312 0.016%
Table 2: Detection and Extraction Performance Across Strategies. While all TASER ablations achieve perfect recall, Full Schema Prompting yields the highest precision (43.43%), F1 score (59.44%), and overall accuracy (66.35%), as well as the lowest total absolute difference (TAD) and unaccounted fraction, underscoring the value of embedding the complete Portfolio schema in the detection prompt. Percentage of unaccounted holdings is out of \$731.7 billion (ground truth). Lower TAD and unaccounted percentages indicate higher dollar‐value fidelity.